OpenStudy (anonymous):

Prove that \[{\tan \phi \over a+\tan^2\phi} = \sin \phi \cos \phi\]

OpenStudy (anonymous):

sorry, a = 1

OpenStudy (mani_jha):

\[1+\tan ^{2}\phi=\sec ^{2}\phi\] so, \[\tan \phi/(1+\tan ^{2}\phi)=\tan \phi/\sec ^{2}\phi\] \[\tan \phi \times \cos ^{2}\phi\] =\[(\sin \phi/\cos \phi)\times \cos ^{2}\phi\] =\[\sin \phi \cos \phi\]

OpenStudy (anonymous):

whys is \[\tan \phi \over \sec^2 \phi\] resulting from the other equation?

OpenStudy (mani_jha):

Because I have replaced \[1+\tan ^{2}\phi \] in the original question with \[\sec ^{2}\phi\] They are equal to each other, isnt it?

OpenStudy (anonymous):

Ok, how does it become \[\tan \phi \times \cos^2 \phi~?\]

OpenStudy (mani_jha):

\[1/\sec ^{2}\phi=\cos ^{2}\phi\] Again replacement of 1/sec^2phi with cos^2phi

sam (.sam.):

(tanx)/(1+tan^(2)x) = sinxcosx Convert the sec^(2)x to its (sin)/(cos) equivalent. \[\huge \frac{(tanx)}{\frac{1}{\cos^2(x)}}\] Remove the single term factors from the expression. tanxcos^(2)(x) Replace tanx with an equivalent expression (sinx)/(cosx) using the fundamental identities. (sinxcos^(2)(x))/(cosx) Reduce the expression (sinxcos(x)^(2))/(cos(x)) by removing a factor of cos(x) from the numerator and denominator =sinxcos(x)

OpenStudy (anonymous):

Oh, Ok.